OpenStudy (anonymous):

A goat is tethered by a 6 metre rope to the outside corner of a shed measuring 4 metres by 5 metres in a grassy field. What area of grass can the goat graze?

OpenStudy (dumbcow):

Since the corner of the shed is a 90 degree angle. That leaves 3/4 of a circle with radius of 6. \[(3/4)*\pi*6^{2} = 27 \pi\] Also , the goat can reach around the corners by 2 on the short sides and 1 on long side. creating 2 90 degree sectors of additional grazing area. \[(1/4)*\pi*2^{2} = \pi\] \[(1/4)*\pi*1^{2} = \pi/4\] \[27 \pi + \pi + \pi/4 = 28.25 \pi\]

OpenStudy (anonymous):

Correct!